# DVLO: Deep Visual-LiDAR Odometry with Local-to-Global Feature Fusion and Bi-Directional Structure Alignment

### v1

@@ 4.2 Implementation Details
Data Preprocessing.   We directly input all LiDAR points without downsampling into our network. Projected pseudo image size is set in line with the range of LiDAR sensor as 64 $\times$ 1800. We pad the camera images to a uniform size of 384 $\times$ 1280. Since there is a large difference in the spatial range captured by camera and LiDAR sensors, we design a fusion mask to indicate which point can be fused with the image.
Parameters.   All experiments are conducted on a single NVIDIA RTX 4090 GPU with PyTorch 1.10.1. We use Adam [16] as our optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$. The initial learning rate is set to 0.001 and exponentially decays every 200,000 steps until it reaches 0.00001. Training batchsize is set as 8. $\alpha^{l}$ in the loss function is set to 1.6, 0.8, 0.4, and 0.2 for four layers. Initial values of learnable parameters $k_{x}$ and $k_{q}$ are set as 0.0 and -2.5, respectively.
@@
Table 2: Comparison with tranditional visual-LiDAR odometry on KITTI 00-10 sequences. Our DVLO is trained on 00-06 sequences. The best results for each sequence are bold, and the second best results are underlined.
@@
Table 3: Comparison with learning-based multi-modal odometry networks on the 09-10 sequences. Our DVLO is trained on 00-06 sequences while other models are trained on 00-08 sequences. The best results for each sequence are bold, and the second best results are underlined.
@@
       Ours       visual+LiDAR       0.85       0.36       0.88       0.48       0.87       0.42
@@ 4.3 Quantitative Results
Comparison with Visual/LiDAR Odometry.   We compare our method with some representative visual odometry (VO) or LiDAR odometry (LO) networks for the comprehensive comparison. Following the settings in [47], we train our model on 00-06 sequences. Quantitative results on the KITTI dataset are listed in Tab. 1, which show that our DVLO outperforms all these works on most sequences. Compared with deep visual odometry, e.g., DFVO[56] and Cho et al.[6], our method’s mean errors $t_{rel}$ and $r_{rel}$ on sequence 07-10 have a 63.4% and a 43.8% decline, respectively. Notably, even though these VO methods are mostly trained on larger data (00-08), our method still outperforms them by a large margin. Compared with deep LiDAR odometry, our DVLO even outperforms the recent SOTA method EfficientLO [46] on most sequences. Compared with EfficientLO, our method has a competitive 0.41 /100m rotation error. Moreover, our mean translation error $t_{rel}$ on testing sequences has a 4.9% decline compared with theirs. The experiment results prove the effectiveness and great potential for our visual-LiDAR fusion design.
Comparison with Traditional Multi-Modal Odometry.   We compare the performance between our method and previous traditional multi-modal odometry works on the whole KITTI sequences (00-10). The results are shown in Tab. 2, which demonstrate that our DVLO outperforms all these works on most sequences. Compared with PL-LOAM[12], our method’s mean translation errors $t_{rel}$ on sequence 00-10 has a 28.7% decline.
Comparison with Learning-based Multi-Modal Odometry.   Because most deep multi-modal fusion odometry methods are trained on the 00-08 sequences and tested on the 09-10 sequences, we also compare the performance between our DVLO and other learning-based multi-modal odometry works on the 09-10 sequences. The results are shown in Tab. 1 and Tab. 3. Notably, even though our model is only trained on the 00-06 sequences, our method still outperforms H-VLO on most sequences where our method has a 47.0% lower $t_{rel}$ and 2.3% lower $r_{rel}$.
Figure 6: Visualization of our designed local clustering-based fusion mechanism within a certain cluster. Red points indicate the 2D positions of cluster centers. And the yellow regions are clustered pixels around each center.
@@ 4.4 Runtime Analysis
Efficiency is another extremely significant factor in real-time SLAM systems. As shown in Tab. 4, we compare the runtime of our DVLO with other multi-modal odometry methods. Since the LiDAR points in KITTI dataset are captured at a 10Hz frequency, previous multi-modal methods [51, 12, 18, 38] rarely satisfy the real-time application requirements (under 100 ms). However, our method has only 98.5 ms inference time, which has the potential for real-time application.
@@
Table 5: Comparison with previous scene flow estimation works on the “val” split of the FlyingThings3D subset. “RGB” and “XYZ” denote the image and point cloud respectively. The best results are bold.
@@ 4.6 Generalization to Scene Flow Estimation Task
It is worth noting that our design can serve as a generic fusion module, which generalizes well to other tasks. Here, we extend our fusion module to the scene flow estimation task. As shown in Tab. 5, our method surpasses all recent SOTA multi-modal scene flow networks on the FlyingThings3D dataset [31] in terms of both 2D and 3D metrics. Our method even consistently outperforms CamLiRAFT [25], which is specially designed for the multi-modal scene flow task. The experiment results demonstrate the strong generalization and universal application capabilities of our method. We also visualize the estimated flow in Fig. 7.
@@ 4.7 Ablation Study
Without Local Fuser.   We remove the Local Fuser module from our network and directly fuse the image and point features using the Global Fuser module. The results in Tab. 6 show that the performance of our model drops significantly without Local Fuser. This demonstrates the importance of the Local Fuser module, since local point-to-pixel correspondences can merge more fine-grained features from different modalities.
Without Global Fuser.   We remove the Global Fuser module and directly use the local fused features for pose estimation. Results in Tab. 6 demonstrate that the limited receptive field of the Local Fuser module can not enable sufficient global information interaction. The global modeling ability of Global Fuser can help to recognize outliers that are harmful to pose regression [27].
@@ 5 Conclusion
In this paper, we propose a novel local-to-global fusion network with bi-directional structure alignment for visual-LiDAR odometry. A clustering-based local fusion module is designed for providing fine-grained multi-modal feature exchange. Furthermore, an adaptive global fusion is designed to achieve global information interaction. Comprehensive experiments show that our DVLO achieves state-of-the-art performance in terms of both accuracy and efficiency. Our fusion module can also serve as a rather generic fusion strategy, which generalizes well onto the multi-modal scene flow estimation task. We leave the generalization on more multi-modal tasks for the future work.

### v3

@@ 4.2 Implementation Details
Data Preprocessing. We directly input all LiDAR points without downsampling. The projected pseudo image size is set in line with the range of the LiDAR sensor as 64 $\times$ 1800. We pad the camera images to a uniform size of 384 $\times$ 1280. Since there is a large spatial range difference between the camera and LiDAR, we design a fusion mask to indicate which point can be fused with the image.
Parameters. Experiments are conducted on an NVIDIA RTX 4090 GPU with PyTorch 1.10.1. We use Adam optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$. The initial learning rate is set to 0.001 and exponentially decays every 200000 steps until 0.00001. Batch size is 8. $\alpha^{l}$ for four layers are 1.6, 0.8, 0.4, and 0.2. Initial values of learnable parameters $k_{x}$ and $k_{q}$ are set as 0.0 and -2.5, respectively.
@@
Table 2: Comparison with traditional visual-LiDAR odometry on KITTI 00-10 sequences. Our DVLO is trained on 00-06 sequences. The best results for each sequence are bold, and the second best results are underlined.
@@
Table 3: Comparison with learning-based multi-modal odometry on KITTI 09-10 sequences. Our DVLO is trained on 00-06 sequences while other models are trained on 00-08 sequences. The best results are bold, and the second best results are underlined.
@@
       Ours       visual+LiDAR       0.85       0.36       0.88       0.46       0.87       0.41
@@ 4.3 Quantitative Results
Comparison with Visual/LiDAR Odometry. We compare our method with some representative visual odometry (VO) or LiDAR odometry (LO) networks for the comprehensive comparison. Following the settings in [55], we train our model on 00-06 sequences. Quantitative results on the KITTI dataset are listed in Tab. 1, which shows that our DVLO outperforms all these works on most sequences. Compared with deep visual odometry, e.g., DFVO[65] and Cho et al.[6], our method’s mean errors $t_{rel}$ and $r_{rel}$ on sequence 07-10 have a 63.4% and a 43.8% decline, respectively. Notably, even though these VO methods are mostly trained on larger data (00-08), our method still outperforms them by a large margin. Compared with deep LiDAR odometry, our DVLO even outperforms the recent SOTA method EfficientLO [54] on most sequences. Compared with EfficientLO, our method has a competitive 0.41 /100m rotation error. Moreover, our mean translation error $t_{rel}$ on testing sequences has a 4.9% decline compared with theirs. The experiment results prove the effectiveness and great potential for our visual-LiDAR fusion design.
Comparison with Traditional Multi-Modal Odometry. We compare the performance between our method and previous traditional multi-modal odometry works on the whole KITTI sequences (00-10). The results are shown in Tab. 2, which demonstrates that our DVLO outperforms all these works on most sequences. Compared with PL-LOAM [16], our method’s mean translation error $t_{rel}$ on sequence 00-10 has a 28.7% decline.
Comparison with Learning-based Multi-Modal Odometry. Because most deep multi-modal fusion odometry methods are trained on the 00-08 sequences and tested on the 09-10 sequences, we also compare the performance between our DVLO and other learning-based multi-modal odometry works on the 09-10 sequences. The results are shown in Tab. 1 and Tab. 3. Notably, even though our model is only trained on the 00-06 sequences, our method still outperforms H-VLO on most sequences where our method has a 47.0% lower $t_{rel}$ and 2.3% lower $r_{rel}$, which proves the superiority of our proposed fusion strategy.
Figure 6: Visualization of our designed local clustering-based fusion mechanism within a certain cluster. Red points indicate the 2D positions of cluster centers. The yellow regions are clustered pixels around each center.
@@ 4.4 Runtime Analysis
Efficiency is another extremely significant factor in real-time SLAM systems. As shown in Tab. 4, we compare the runtime of our DVLO with other multi-modal odometry methods. Since the LiDAR points in the KITTI dataset are captured at a 10Hz frequency, previous multi-modal methods [57, 16, 23, 47] rarely satisfy the real-time application requirements (under 100 ms). However, our method has only 98.5 ms inference time, which has the potential for real-time application.
@@
Table 5: Comparison with previous scene flow estimation works on the “val” split of the FlyingThings3D subset [39]. “RGB” and “XYZ” denote the image and point cloud respectively. The best results are bold.
@@ 4.6 Generalization to Scene Flow Estimation Task
It is worth noting that our design can serve as a generic fusion module, which generalizes well to other tasks. Here, we extend our fusion module to the scene flow estimation task [36, 29, 32, 20, 12, 11, 68]. As shown in Tab. 5, our method surpasses all recent SOTA multi-modal scene flow networks on the FlyingThings3D dataset [39] in terms of both 2D and 3D metrics. Our method even consistently outperforms CamLiRAFT [29], which is specially designed for the multi-modal scene flow task. The experiment results demonstrate the strong generalization and universal application capabilities of our method. We also visualize the estimated flow in Fig. 7.
@@ 4.7 Ablation Study
Without Local Fuser. We remove the Local Fuser module from our network and directly fuse the image and point features using the Global Fuser module. The results in Tab. 6 show that the performance of our model drops significantly without Local Fuser. This demonstrates the importance of the Local Fuser module since local point-to-pixel correspondences can merge more fine-grained features from different modalities.
Without Global Fuser. We remove the Global Fuser module and directly use the local fused features for pose estimation. Results in Tab. 6 demonstrate that the limited receptive field of the Local Fuser module can not enable sufficient global information interaction, which leads to about 13.4% higher $t_{rel}$ and 14.6% higher $r_{rel}$. The global modeling ability of Global Fuser can help to recognize outliers that are harmful to pose regression [31].
@@ 5 Conclusion
In this paper, we propose a novel local-to-global fusion network with the bi-directional structure alignment for visual-LiDAR odometry. A clustering-based local fusion module is designed to provide fine-grained multi-modal feature exchange. Furthermore, an adaptive global fusion is designed to achieve global information interaction. Comprehensive experiments show that our DVLO achieves state-of-the-art performance in terms of both accuracy and efficiency. Our fusion module can also serve as a rather generic fusion strategy, which generalizes well onto the multi-modal scene flow estimation task. We leave the generalization experiments on more multi-modal tasks for future work.
Acknowledgements
This work was supported in part by the Natural Science Foundation of China under Grant 62225309, 62073222, 62361166632, and U21A20480.